A stapler works by shearing one staple off of a row of staples. Assume that the staple has a bonded thickness of 0.05 alon
Answer #1

The bonded area of stapler pin, A = 2 x (0.3 X 0.05) + 0.5 x 0.05 = 0.055 in2

(a) Let us assume Force F = 1.2 #

Shear stress = F/A = 1.2 / 0.055 = 21.81 psi

(b) Let us assume Force F = 4.8 #

Shear stress = F / A = 4.8/0.055 = 87.27 psi